you can skip Q 21

21.A company estimates an NPV of a project under three different set of assumptions (Bear, Base, Bull) to evaluate forecasting risk; management agrees to undertake the project if the weighted average NPV for the three different scenarios (Bear, Base, Bull) is positive. Based on the scenario analysis performed, the company will pursue the proiect. Evaluate the underlined words in italics. True or False? Scenerio Bear Base Bull NPV Probability 30.00% 35| 50.00% 65 20.00% ru b. False Today is January 1t, 2019 (T-0). You take out a 6 year fully amortizing auto loan of $24,000. Payments are made at the end of each calendar month. The loan has a fixed annual rate of 4.0% (or 4.0%/12 per month). 22. Calculate the monthly payment on the auto loan. If you were to pay off the balance of the loan at the end of the 2nd month (immediately after making the second monthly payment), the amount of money you would owe the auto dealership is closest to: a. $23,239 b. $23,371 c. $23,408 d. $23,666 e. $23,705 23. Ignore the monthly payment amount you calculated in the previous problem. Assume you decide to make monthly payments of $503.50 instead. Approximately how many months sooner would you pay off the auto loan? b. 20 c. 22 e. 48 24. During lecture we illustrated how the profitability of different types of companies can change as the economy moves through a business cycle of 7-10 years. A company like Molson-Coors Brewing Company (which trades under the ticker TAP) would best be described asa When economic conditions deteriorate (e.g. during the housing crisis), TAPs operating margin would most likely a. non-cyclical company b. cyclical compan)y c. non-cyclical company d. cyclical company e. non-cyclical company increase significantly decrease significantly decrease significantly remain the samee remain the same 25. A creditor is willing to offer you $5,000 today in the form of a pure discount loan, an interest only loan, or an amortizing loan. Each loan is for an 8 year period where you will pay interest one time per year (if applicable). While you are willing to pay the creditor annual interest, you do not want to return any of the $5,000 in principle for 8 years. You would also prefer to pay as little interest over the lifetime of the loan as possible. The best loan (s) for you to choose (if any) would be the: a. The pure discount loan b. The interest only loan c. The amortizing loan d. None of the loans satisfy your constraints e. Any of loans would work as they are all for the same amount of money

5. Assume you are a sell-side analyst and you cover the widget industry. In 2018, Big Widget Company decided to double manufacturing capacity by issuing debt in order to capitalize on the growing demand for widgets. Assume that the company was able to realize scale economies through increased buying power. Any benefit from these scale economies was offset by borrowing money at a much higher interest rate than existing long term debt. Using 5-Stage DuPont ROE formula, which of the following would be least likely:

a. In 2018, NI / EBT (or the tax burden ratio) was almost the same as it was in 2017

b. In 2018, EBT / EBIT (or the interest burden ratio) was higher than it was in 2017

c. In 2018, EBIT / Sales (or the operating margin) was higher as it was in 2017

d. In 2018, Sales / TA (or total asset turnover ratio) was higher than it was in 2017

e. In 2018, ROE was almost the same as it was in 2017
Answer #1

As per rules I am answering the first 4 subparts of the question

21: TRUE

Expected NPV= Sum of probability*NPV

= -100*0.3 + 35*0.5 + 65*0.2

= 0.5

Since it is positive, the company will pursue the project

22: c

Using financial calculator

Input: PV = -24000 , N= 6*12= 72, I/Y= 4%/12 = 0.003333

PMT = Monthly payment = $375.48

The amortization table is

Month

StartingBalance

Interest

Principal

EndingBalance

TotalInterest

1

$24,000.00

$80.00

$295.48

$23,704.52

$80.00

2

$23,704.52

$79.02

$296.47

$23,408.05

$159.02

Hence amount remaining = 23408.05

WORKINGS

File Home InsertDraw Page Layout FormulasData Review View Tell me what you want to do Share 10 A A b Wrap Text General Paste e.0 .00 Conditional Format as Cell Insert Delete Format . Sort & Find & E- Merge & Center . $ , % , Formatting Table Styles- Filter Select Clipboard Font Alignment Number Styles Cells Editing 4 5 Loan Amount Interest Rate Term in Years Monthly Payment PMT(B6/12,(C6 12),-A6) Month Principal StartingBalance FA6 Interest EndingBalance Totallnterest -B9-D9 -C9 10 2 11 3 12 4 13 5 14 6 15 7 16 8 B10 (BS6/12) DS6-C10 -81 1·(BS6/12) =DS6-C11 B12 (BS6/12) DS6-C12 B13 (BS6/12) DS6-C13 B14 (BS6/12) DS6-C14 B15 (BS6/12) DS6-C15 B16* (BS6/12) DS6-C16 B10-D10 B11-D11 -B12-D12 -B13-D13 B14-D14 B15-D15 B16-D16 =E10 -E11 -E12 -F10+C11 11+C 12 =F12+C13 F13+C14 =F14+C15 -F15+C16 E13 -E14 -E15 Quarterly Semiannual Monthly Annual Ready + 100% ENG 6:20 AM

23:  B

Using financial calculator
Input: PV= -24000, PMT= 503.5, I/Y = 4%/12 =  0.003333

Find N= 52 months

Difference = 6*12- 52 = 20 months

24: c

It is a non cyclical company. When the economy deteriorates, people will spend lesser and so its sales and operating margin would decline.